# Simulation as Supervision: Mechanistic Pretraining for Scientific Discovery

### v1

@@ 1 Abstract
Scientific modeling faces a core limitation: mechanistic models offer interpretability but collapse under real-world complexity, while machine learning models are flexible but require large labeled datasets, cannot infer unobservable quantities, and operate as black boxes. We introduce Simulation-Grounded Neural Networks (SGNNs), a general framework that uses mechanistic simulations as training data for neural networks. SGNNs are pretrained on synthetic corpora spanning diverse model structures, parameter regimes, stochasticity, and observational artifacts. We evaluated SGNNs across scientific disciplines and modeling tasks, and found that SGNNs achieved state-of-the-art results across settings: for prediction tasks, they nearly tripled COVID-19 forecasting skill versus CDC baselines, reduced chemical yield prediction error by one-third, and maintained accuracy in ecological forecasting where task-specific models failed. For inference tasks, SGNNs also accurately classified the source of information spread in simulated social networks and enabled supervised learning for unobservable targets, such as estimating COVID-19 transmissibility more accurately than traditional methods even in early outbreaks. Finally, SGNNs enable back-to-simulation attribution, a new form of mechanistic interpretability. Given a real-world input, SGNNs retrieve simulations based on what the model has learned to see as most similar, revealing which underlying dynamics the model believes are active. This provides process-level insight—what the model thinks is happening—not just which features mattered. SGNNs unify scientific theory with deep learning flexibility and unlock a new modeling paradigm—transforming simulations from rigid, post hoc tools into flexible sources of supervision, enabling robust, interpretable inference even when ground truth is missing.
@@ 2 Introduction
Scientific modeling has long been divided between two approaches: theory-driven models, which encode domain knowledge through mechanistic equations, and data-driven models, which learn flexible patterns from large datasets. Each has fundamental limitations. Mechanistic models are interpretable and grounded in scientific theory, but often break down in practice due to noise, unobserved variables, or incorrect structural assumptions. Neural networks, by contrast, can capture complex relationships but require large labeled datasets, lack scientific grounding, and often learn spurious patterns that fail when applied to new data—limiting their utility for scientific reasoning.
This tension is especially acute in fields like epidemiology, ecology, chemistry, and social science, where real-world data are noisy, sparse, and biased—and where key variables like transmission rates or carrying capacities are unobservable. The result is a core bottleneck: we lack models that are both scientifically grounded and capable of robust generalization from imperfect data.
Several approaches have been proposed to bridge this divide. Physics-Informed Neural Networks (PINNs) incorporate governing equations directly into model optimization, promoting physical consistency in learned dynamics [1, 2]. These methods perform well when equations are known and all relevant variables are observable. However, in many scientific settings, mechanistic models are incomplete—omitting stochasticity, feedbacks, or spatial effects—and enforcing them during training can introduce bias, especially under regime shifts. Other approaches use simulation-trained neural networks for surrogate modeling or task-specific prediction, often achieving strong results in constrained settings [3]. Yet, many such methods depend on a single simulator or fixed structure with hand-tuned parameters, limiting generalization.
@@
We introduce Simulation-Grounded Neural Networks (SGNNs), a general framework that uses mechanistic simulations to train neural networks. SGNNs treat mechanistic models as generators of diverse, labeled training data—spanning alternative model structures, parameter regimes, stochasticity, and realistic observational artifacts. This allows the neural network to internalize scientific structure without requiring it to be hard-coded into the architecture. SGNNs therefore unify the flexibility of deep learning with the theory-driven approach of mechanistic science by using simulation as supervision.
@@
SGNNs are aligned with recent work demonstrating that large neural networks can acquire general-purpose inference strategies through synthetic supervision [4, 5, 6]. These models show that abstract generative processes can produce strong reasoning priors. SGNNs build on this foundation but differ fundamentally in being grounded in domain-specific scientific simulations—yielding models that are not only performant, but mechanistically interpretable and aligned with scientific theory. In this sense, SGNNs act as scientific inference engines: trained on simulations, and capable of reasoning under real-world uncertainty.
@@
This approach directly addresses a core limitation of machine learning in science: the inability to infer unobservable quantities. Scientific parameters like the basic reproduction number in an infectious disease outbreak (the expected number of new infections from a single case) or ecological carrying capacity can’t be measured directly and therefore lack labeled data. Traditional methods estimate them by fitting mechanistic models to observed trajectories, but this process is fragile—sensitive to noise, delays, and structural mismatch. SGNNs bypass these issues by training on synthetic data where ground truth is known, enabling robust supervised learning for latent scientific quantities.
@@
SGNNs also enable a new form of interpretability we call back-to-simulation attribution. After training, the model embeds real-world inputs into its internal representation space and retrieves the most similar simulations from its synthetic training data. These retrieved examples reflect the mechanistic regimes the model associates with the input—revealing what kinds of system dynamics it believes are driving the observed data. Because the retrieved simulations have known ground-truth parameters, users can directly inspect the model’s implicit beliefs. For example, an outbreak with few reported cases might be matched to simulations with high asymptomatic spread and delayed reporting—suggesting the model interprets it as a largely hidden early-stage epidemic. Unlike post hoc feature attribution methods such as SHAP [7] or LIME [8], back-to-simulation attribution provides process-level explanations grounded in mechanistic structure—not just which input features mattered, but which scientific mechanisms the model thinks are at play.
@@
We demonstrate SGNNs across tasks in epidemiology, ecology, chemistry, and social science—showing strong generalization in forecasting, regression, classification, and parameter inference. SGNNs nearly triple the forecasting skill of the CDC Forecast Hub’s median model on early COVID-19, despite never seeing real COVID-19 data. Additionally, SGNNs trained only on human-to-human disease simulations outperform specialized models on dengue (a vector-borne disease) by 24%, demonstrating the robustness of the approach. In chemical reaction yield prediction, SGNNs reduce residual variance by one-third with only three minutes of pretraining, and in ecological forecasting, SGNNs retain high accuracy in high-dimensional settings where task-specific neural nets fail completely.
@@
Together, these results position SGNNs as a general modeling paradigm for scientific machine learning—one that transforms mechanistic simulations into flexible supervisory signals, enabling robust, interpretable, and data-efficient learning across disciplines. [^1] Code and data available at https://github.com/carsondudley1/SGNNs. This includes full model training and synthetic data generation pipelines for all domains studied.
@@ 3.1 Conceptual Framework
SGNNs train neural networks on synthetic data generated by mechanistic simulators. Let $S(\theta)$ denote a stochastic simulator parameterized by scientific parameters $\theta$, which generates trajectories or time series $x_{i}=S(\theta_{i})$. To capture a wide range of system behaviors, we sample $\theta_{i}\sim P(\theta)$ from a task-specific distribution that spans multiple model classes, parameter regimes, and observational settings.
Each simulation yields one or more training examples $x_{i}$ paired with targets $y_{i}=T(\theta_{i})$, where $T(\cdot)$ defines the labeling function—e.g., future trajectories, final states, or latent scientific quantities. Importantly, simulations are passed through an observation model that adds realistic artifacts such as noise, censoring, delays, or reporting bias. The result is a synthetic dataset $\mathcal{D}_{\text{syn}}=\{(x_{i},y_{i})\}$ that reflects both mechanistic diversity and realistic surveillance scenarios.
A neural network $f_{\phi}$ is then trained via supervised learning to predict $y_{i}$ from $x_{i}$. Because each input is paired with exact labels—including latent scientific quantities—and the training set spans a wide range of mechanistic regimes and observational conditions, the model learns to internalize structure that generalizes beyond any single model or parameter setting. SGNNs are architecture-agnostic: any neural network with sufficient capacity may serve as $f_{\phi}$. In practice, we use domain-specific backbones such as temporal transformers, graph networks, or hybrid models (Appendix B).
@@ 3.2 Key Capabilities
  • Mechanistic interpretability. Through back-to-simulation attribution, SGNNs identify which simulated regimes they internally associate with real-world inputs—offering insight into the underlying dynamics the model believes are active, and enabling process-level explanations.
@@ 4.2.1 SGNNs Outperform CDC COVID-19 Forecasting Models
The SGNN was pretrained entirely on large-scale mechanistic simulations, which incorporated a wide range of epidemic dynamics—including multiple transmission pathways, contact structures, latent periods, stochastic super-spreading, policy interventions, seasonality, underreporting, and surveillance delays (Appendix A). No real-world or COVID-specific data were used in pretraining.
We assessed performance using forecasting skill, defined as the percentage improvement in mean absolute error (MAE) relative to a naive baseline that simply predicts the most recent value. This is a stringent metric in infectious disease forecasting, where trends are often flat or noisy. Indeed, the mean skill across all models in the CDC Forecast Hub was negative (–1.0) [10], reflecting widespread failure to outperform the naive baseline. For comparison, we benchmarked the SGNN against both the median skill (13.0) and best model’s skill (34.0) of Forecast Hub models.
The SGNN outperformed all baselines (Figure 2). Compared to Chronos [11]—a state-of-the-art foundation model pretrained on generic time series—the SGNN achieved an over 12-fold improvement in skill, underscoring the importance of mechanistic grounding. The SGNN also nearly tripled the Forecast Hub’s median skill and exceeded its best model performance—despite being trained entirely on synthetic data and applied out-of-the-box to real COVID-19. Finally, the SGNN outperformed a physics-informed neural network (PINN) baseline based on SEIR equations (skill = –112), which performed poorly in our tests in the presence of noise and data sparsity, consistent with prior observations [2].
@@ 4.2.2 SGNNs Forecast Accurately Despite Mechanistic Misspecification
Mechanistic models—whether traditional compartmental models or hybrid approaches like PINNs—rely on structural assumptions that rarely hold in practice. Real-world systems often deviate from these assumptions, leading to brittle forecasts when the model form is even slightly wrong. This challenge is especially acute when key mechanisms (e.g., transmission routes or feedback loops) are misrepresented or omitted. To assess SGNNs under such conditions, we tested their forecasting performance on a domain with fundamentally different mechanisms from those seen during pretraining.
We evaluated forecasting performance on real-world dengue outbreaks in Brazil [12]. Dengue transmission involves mosquito vectors and mechanisms that are distinct from the human-to-human transmission used in all SGNN pretraining simulations. Importantly, SGNNs were evaluated in a fully zero-shot setting: the model had no exposure to dengue data or vector-borne transmission processes during training.
To benchmark performance, we compared SGNNs against two baselines. The first was a PINN-style model trained directly on the dengue data using standard SEIR equations—imposing a rigid, incorrect mechanistic prior [2]. The second was a purely statistical model based on exponential smoothing (ETS), which had access to the dengue data but incorporated no mechanistic assumptions.
@@
These findings underscore a defining strength of SGNNs: by training on a broad ensemble of mechanistic simulations spanning multiple model structures and parameter regimes, SGNNs learn flexible, scientifically grounded representations that generalize across domains—even when key structural elements differ from those seen in training. Crucially, this result suggests that SGNNs do not merely memorize surface patterns, but instead internalize scientific principles that extend across mechanistic boundaries. This result suggests that SGNNs may be learning abstract representations that capture core features of disease spread—such as delays, feedback, and depletion—that hold across different mechanistic forms.
@@ 4.3 Realistic, Diverse Mechanistic Grounding is Essential for Generalization
First, we replaced our simulation pipeline with a neural network trained directly on COVID-19 case data. This neural model acted as a simulator: it generated synthetic outbreaks by recursively forecasting forward from real historical contexts. We then trained a model with the same architecture as our SGNN directly on this simulated data. Despite the model having access to high amounts of simulated data, it was significantly outperformed by the SGNN trained on mechanistic simulations, yielding over twice the error (Figure 2e). This experiment demonstrates that increasing data alone is insufficient—mechanistic grounding is critical for stable and generalizable learning.
@@ 4.4.1 SGNNs Forecast Multispecies Ecological Dynamics Where Neural Nets Fail
SGNNs achieved strong forecasting skill on both lynx and hare populations. As shown in Figure 3A (left), SGNNs outperformed statistical and mechanistic baselines, with forecasting skill of 0.48 for lynx and 0.40 for hare—substantially higher than the 0.2–0.32 range for all baselines. These results highlight SGNNs ability to generalize to cyclic ecological systems from limited data.
On higher-dimensional forecasting, SGNNs maintained robust forecasting skill as the number of butterfly species increased from 4 to 16 (Figure 3A, right), outperforming task-specific neural networks, which failed entirely in high dimensions. By 16 species, SGNN retained positive forecasting skill above 0.19, while baseline neural nets fell below zero. This illustrates SGNN’s ability to exploit shared structure learned during pretraining, enabling effective high-dimensional forecasting where standard approaches break down.
@@ 4.4.2 SGNNs Boost Chemical Yield Prediction with Minimal Pretraining
State-of-the-art data-driven models trained directly on this dataset leave 15% of the variance unexplained on held-out test data [18]. SGNNs, pretrained for only a few minutes on high-fidelity chemical reaction network simulations generated via a hybrid simulation pipeline (Appendix A), reduced residual variance to 10%—a one-third reduction. This improvement reflects SGNN’s ability to internalize mechanistic dependencies such as catalyst-substrate incompatibilities, reaction failures, and heteroscedastic noise that purely empirical models fail to capture from limited data alone. A major weakness of purely data-driven models was their inability to distinguish failed reactions, which often led to overconfident but inaccurate predictions in low-yield regimes (Figure 3B, right). By explicitly encoding failure probabilities into our synthetic simulator, SGNNs learned to recognize reactions likely to fail—resulting in sharper, more calibrated predictions across the full yield distribution. As shown in Figure 3B, SGNNs yield more accurate predictions with lower residual variance and better alignment with ground-truth yields.

### v2

@@ 1 Abstract
Scientific modeling faces a tradeoff: mechanistic models provide scientific grounding but struggle with real-world complexity, while machine learning models achieve strong predictive performance but require large labeled datasets and are not interpretable. We introduce Simulation-Grounded Neural Networks (SGNNs), which use mechanistic simulations as training data for neural networks. SGNNs are pretrained on synthetic corpora spanning diverse model structures, parameter regimes, stochasticity, and observational artifacts. Simulation-grounded learning has been applied in multiple domains (e.g., surrogate models in physics, forecasting in epidemiology). We provide a unified framework for simulation-grounded learning and evaluated SGNNs across scientific disciplines and modeling tasks. We found that SGNNs were successful across domains: for prediction tasks, they nearly tripled COVID-19 forecasting skill versus CDC baselines, reduced chemical yield prediction error by one-third, and maintained accuracy in ecological forecasting where task-specific models failed. For inference tasks, SGNNs also accurately classified the source of information spread in simulated social networks and enabled supervised learning for unobservable targets, such as estimating COVID-19 transmissibility more accurately than traditional methods even in early outbreaks. Finally, SGNNs enable back-to-simulation attribution, a form of mechanistic interpretability. Back-to-simulation attribution matches real-world observations to the training simulations the model considers most similar, identifying which mechanistic processes the model believes best explain the observed data. By providing a unified framework for simulation-grounded learning, we establish when and how mechanistic simulations can serve as effective training data for robust, interpretable scientific inference.
@@ 2 Introduction
Scientific modeling is often divided between two approaches: theory-driven models, which encode domain knowledge into mathematical equations, and data-driven models, which learn flexible patterns from large datasets. Each has fundamental limitations. Mechanistic models are interpretable and grounded in scientific theory, but often break down in practice due to noise, unobserved variables, or incorrect structural assumptions. Neural networks can capture complex relationships but require large labeled datasets, lack scientific grounding, and often learn spurious patterns that fail when applied to new data. We lack models that are both scientifically grounded and capable of robust generalization from imperfect data.
Recent work has sought to bridge this divide by incorporating mechanistic knowledge into machine learning. These approaches can be understood through the three core components of machine learning problems: the data, the model architecture, and the optimization procedure. Physics-Informed Neural Networks (PINNs) incorporate governing equations into the loss function, constraining optimization to satisfy known physical laws [1, 2]. Neural ODEs embed continuous-time dynamics directly into the architecture, parameterizing derivative functions that are solved during forward passes [3]. Universal Differential Equations (UDEs) combine both approaches, using neural networks to learn unknown terms in otherwise known equations while training on trajectory data [4]. These methods perform well when equations are known and all relevant variables are observable, but can introduce bias when mechanistic models are incomplete or when systems undergo regime shifts.
Another approach is to incorporate mechanistic knowledge into the training data itself: using simulations from mechanistic models to generate synthetic datasets for neural network training. This strategy has emerged across disciplines, driven by different motivations. In physics, simulation-trained networks can serve as fast surrogates for expensive simulators, enabling real-time inference where full simulations are computationally prohibitive [5]. In epidemiology and social science, simulations address data scarcity and enable learning of unobservable quantities [6]. Despite strong empirical results in specific applications, these efforts lack a unified framework for understanding when and why simulation-grounded learning succeeds.
@@
We introduce Simulation-Grounded Neural Networks (SGNNs) as a general framework for training neural networks on mechanistic simulations. While simulation-grounded learning has been applied in isolated contexts, we provide a unified treatment and evaluation across domains. SGNNs are pretrained on synthetic corpora spanning diverse model structures, parameter regimes, stochasticity, and realistic observational artifacts. This approach treats mechanistic models as generators of labeled training data, allowing neural networks to internalize scientific structure without requiring it to be incorporated into architectures or loss functions. This gives a flexible framework compatible with any architecture or optimization method.
@@
SGNNs are aligned with recent work demonstrating that large neural networks can acquire general-purpose inference strategies through synthetic supervision [7, 8, 9]. These models show that abstract generative processes can produce strong reasoning priors. SGNNs build on this foundation but differ fundamentally in being grounded in domain-specific scientific simulations. This yields models that are not only performant, but mechanistically interpretable and aligned with scientific theory. In this sense, SGNNs act as scientific inference engines: trained on simulations, and capable of reasoning under real-world uncertainty.
@@
This approach directly addresses a core limitation of machine learning in science: the inability to infer unobservable quantities. Scientific parameters like the basic reproduction number in an infectious disease outbreak or ecological carrying capacity can’t be measured directly. Traditional methods estimate them by fitting mechanistic models to observed trajectories, a process sensitive to noise, delays, and structural mismatch. Simulation-based inference (SBI) addresses this by using simulation-grounded training to learn posterior distributions over parameters [10, 11, 12], providing a principled approach to Bayesian parameter inference. However, full posterior inference is not always necessary: simulation-grounded learning can also enable simpler approaches such as direct quantile regression on parameters, which can achieve accurate inference with reduced computational complexity.
@@
SGNNs also enable back-to-simulation attribution, a form of mechanistic interpretability. After training, the model retrieves training simulations that it considers most similar to real-world inputs, revealing which mechanistic regimes the model associates with observed data. Because retrieved simulations have known ground-truth parameters, users can directly inspect the model’s implicit beliefs about underlying processes. Unlike post hoc feature attribution methods such as SHAP [13] or LIME [14], back-to-simulation attribution provides process-level explanations: not just which input features mattered, but which scientific mechanisms the model believes are active.
@@
We demonstrate SGNNs across tasks in epidemiology, ecology, chemistry, and social science, showing strong generalization in forecasting, regression, classification, and parameter inference. SGNNs nearly triple the median forecasting skill of the CDC Forecast Hub on early COVID-19 data, despite never seeing real-world disease data. Additionally, SGNNs trained only on human-to-human disease simulations outperform specialized models on dengue (a vector-borne disease) by 24%, demonstrating the robustness of the approach. In chemical reaction yield prediction, SGNNs reduce residual variance by one-third, and in ecological forecasting, SGNNs retain high accuracy in high-dimensional settings where task-specific neural nets fail completely.
@@
These results establish SGNNs as a general framework for scientific machine learning, demonstrating when and how mechanistic simulations can serve as effective training data for robust, interpretable inference across disciplines. [^1] Code and data available at https://github.com/carsondudley1/SGNNs.
@@ 3.1 Conceptual Framework
SGNNs train neural networks on synthetic data generated by mechanistic simulators. Let $S(\theta)$ denote a stochastic simulator parameterized by $\theta$, which generates trajectories or time series $x_{i}=S(\theta_{i})$. To capture a wide range of system behaviors, we sample $\theta_{i}\sim P(\theta)$ from a task-specific distribution that spans multiple model classes, parameter regimes, and observational settings.
Each simulation yields one or more training examples $x_{i}$ paired with targets $y_{i}=T(\theta_{i})$, where $T(\cdot)$ defines the labeling function (e.g., future trajectories, final states, or latent scientific quantities). Importantly, simulations are passed through an observation model that adds realistic artifacts such as noise, censoring, delays, or reporting bias. The result is a synthetic dataset $\mathcal{D}_{\text{syn}}=\{(x_{i},y_{i})\}$ that reflects both mechanistic diversity and realistic surveillance scenarios.
A neural network $f_{\phi}$ is then trained via supervised learning to predict $y_{i}$ from $x_{i}$. Because each input is paired with exact labels, including latent scientific quantities, and the training set spans a wide range of mechanistic regimes and observational conditions, the model learns to internalize structure that generalizes beyond any single model or parameter setting. SGNNs are architecture-agnostic: any neural network with sufficient capacity may serve as $f_{\phi}$. In practice, we use domain-specific backbones such as temporal transformers, graph networks, or hybrid models (Appendix B).
@@ 3.2 Key Capabilities
  • Mechanistic interpretability. Through back-to-simulation attribution, SGNNs identify which simulated regimes they internally associate with real-world inputs, offering insight into the underlying dynamics the model believes are active, and enabling process-level explanations.
@@ 4.2.1 SGNNs Outperform CDC COVID-19 Forecasting Models
The SGNN was pretrained entirely on large-scale mechanistic simulations, which incorporated a wide range of epidemic dynamics, including stochastic super-spreading, policy interventions, seasonality, underreporting, and surveillance delays (Appendix A). No real-world or COVID-specific data were used in pretraining.
We assessed performance using forecasting skill, defined as the percentage improvement in mean absolute error (MAE) relative to a naive baseline that predicts the most recent value:
$\text{Skill}=100\times\left(1-\frac{\text{MAE}_{\text{model}}}{\text{MAE}_{\text{baseline}}}\right)$ (1)
where the baseline prediction is $\hat{y}_{t+h}=y_{t}$ for all forecast horizons $h$. This is a stringent metric in infectious disease forecasting. The mean skill across all models in the CDC Forecast Hub was negative (–1.0) [16]. For comparison, we benchmarked the SGNN against both the median skill (13.0) and best model’s skill (34.0) of Forecast Hub models.
The SGNN outperformed the baselines (Figure 2). Compared to Chronos [17], a state-of-the-art foundation model pretrained on generic time series, the SGNN achieved a roughly 12-fold improvement in skill, underscoring the importance of mechanistic grounding. The SGNN also nearly tripled the Forecast Hub’s median skill and slightly exceeded its best model performance. Finally, the SGNN outperformed a physics-informed neural network (PINN) baseline based on SEIR equations (skill = –112), which performed poorly in our tests in the presence of noise and data sparsity, consistent with prior observations [2]. This demonstrates that incorporating domain knowledge into the data rather than constraining the output of a neural network to match specific equations may be a better strategy in the presence of large uncertainty and minimal real-world data.
@@ 4.2.2 SGNNs Forecast Accurately Despite Mechanistic Misspecification
Mechanistic models—whether traditional compartmental models or hybrid approaches like PINNs—rely on structural assumptions that may not hold in practice. Real-world systems often deviate from these assumptions, leading to brittle forecasts when the model form is even slightly wrong. This challenge is especially acute when key mechanisms (e.g., transmission routes or feedback loops) are misrepresented or omitted. To assess SGNNs under such conditions, we tested their forecasting performance with different mechanisms from those seen during pretraining.
We evaluated forecasting performance on real-world dengue outbreaks in Brazil [18]. Dengue transmission involves mosquito vectors and mechanisms that are distinct from the human-to-human transmission used in SGNN pretraining simulations. Importantly, SGNNs were evaluated in a fully zero-shot setting: the model had no exposure to dengue data or vector-borne transmission processes during training.
To benchmark performance, we compared SGNNs against two baselines. The first was a PINN-style model trained directly on the dengue data using standard SEIR equations, imposing a rigid, incorrect mechanistic prior [2]. The second was a purely statistical model based on exponential smoothing (ETS), which had access to the dengue data but incorporated no mechanistic assumptions.
@@
These findings underscore a strength of SGNNs: by training on a broad ensemble of mechanistic simulations spanning multiple model structures and parameter regimes, SGNNs learn flexible, scientifically grounded representations that can generalize when structural elements differ from those seen in training. This result suggests that SGNNs do not merely memorize surface patterns, but instead internalize scientific principles that extend across mechanistic boundaries. SGNNs may be learning abstract representations that capture core features of disease spread—such as delays, feedback, and depletion—that hold across different mechanistic forms.
@@ 4.3 Realistic, Diverse Mechanistic Grounding is Essential for Generalization
First, we replaced our simulation pipeline with a neural network trained directly on COVID-19 case data. This neural model acted as a simulator: it generated synthetic outbreaks by recursively forecasting forward from real historical contexts. We then trained a model with the same architecture as our SGNN directly on this simulated data. Despite the model having access to high amounts of simulated data, it was significantly outperformed by the SGNN trained on mechanistic simulations, yielding over twice the error (Figure 2e). This experiment demonstrates that increasing data alone is insufficient. Mechanistic grounding is critical for stable and generalizable learning.
@@ 4.4.1 SGNNs Forecast Multispecies Ecological Dynamics Where Neural Nets Fail
SGNNs achieved strong forecasting skill on both lynx and hare populations. As shown in Figure 3A (left), SGNNs outperformed statistical and mechanistic baselines, with forecasting skill of 48.0 for lynx and 40.2 for hare, higher than the 20.1-32.4 range for all baselines. These results highlight SGNNs ability to generalize to cyclic ecological systems from limited data.
On higher-dimensional forecasting, SGNNs maintained robust forecasting skill as the number of butterfly species increased from 4 to 16 (Figure 3A, right), outperforming task-specific neural networks, which failed entirely in high dimensions. By 16 species, SGNN retained positive forecasting skill above 19.0, while baseline neural nets fell below zero. SGNNs can enable high-dimensional forecasting where standard approaches break down.
@@ 4.4.2 SGNNs Boost Chemical Yield Prediction with Minimal Pretraining
State-of-the-art data-driven models trained directly on this dataset leave 15% of the variance unexplained on held-out test data [24]. SGNNs, pretrained for only a few minutes on high-fidelity chemical reaction network simulations generated via a hybrid simulation pipeline (Appendix A), reduced residual variance to 10%, a one-third reduction. This improvement reflects SGNN’s ability to internalize mechanistic dependencies such as catalyst-substrate incompatibilities, reaction failures, and heteroscedastic noise that purely empirical models fail to capture from limited data alone. A major weakness of purely data-driven models was their inability to distinguish failed reactions, which often led to overconfident but inaccurate predictions in low-yield regimes (Figure 3B, right). By explicitly encoding failure probabilities into our synthetic simulator, SGNNs learned to recognize reactions likely to fail, leading to sharper, more calibrated predictions across the full yield distribution. As shown in Figure 3B, SGNNs yield more accurate predictions with lower residual variance and better alignment with ground-truth yields.